# A Subgrid Model of Electron-Scale Tokamak Turbulence

### v1

@@ Abstract
A subgrid electron-temperature-gradient (ETG) model is demonstrated here which averages local electron-scale turbulence from the GENE code over intermediate scales in space and time to include in global ion-temperature-gradient (ITG) GEM simulations. This approach results in ion-scale equations which include the electron heat transport from ETG turbulence and the effects of electron-scale turbulence on the ion scale. Flux-tube ETG Cyclone Base Case simulations are carried out using GENE at different radial locations and a kinetic form of the flux is added to ion-scale global GEM simulations as a source term. Analytic radial profiles of ETG heat flux are constructed and compared to flux-tube results at multiple radial locations. Different ratios of ITG to ETG turbulent heat flux levels are considered and the results of capturing ETG heat transport in global GEM simulations are discussed. Possibilities for coupling of the ETG streamer potential and intermediate-scale zonal flows found in ETG simulations to the ion scale are further discussed.
@@ I Introduction
The role of ETG turbulence has been studied in various tokamak scenariosGurchenko and Gusakov (2010); Chapman et al. (2015); Ren et al. (2017); Ryter et al. (2019); Kiefer et al. (2021) and is particularly important in cases of suppressed ion-scale turbulence.Ren et al. (2024) Specifically for ITER, alpha-particle and electron cyclotron heating effects are expected to drive meaningful ETG turbulence levels, leading to important multiscale dynamics when ITG turbulence is marginal. Holland, Howard, and Grierson (2017); Howard et al. (2017) To better understand interactions between the disparate scales, local multiscale simulations of core ITG and ETG turbulence have garnered much interest.Jenko (2004); Maeyama et al. (2015); Howard et al. (2016); Holland, Howard, and Grierson (2017); Howard et al. (2017); Bonanomi et al. (2018); Howard et al. (2021) In general, turbulent spectra are distinctly scale-separated, and cross-scale interactions lead to changes in steady-state transport levels which can better predict experimental losses for both species. A recent overview of multiscale simulation results can be found in Ref. Maeyama et al., 2024.
As multiscale simulations require resolving electron gyroradius scales, the sizes of simulation domains become limited. Consequently, reduced modeling of electron-scale turbulence is valuable for whole-device modeling efforts in future burning plasma experiments. Previous theoretical work has considered the importance of cross-scale interactions,Itoh et al. (2002); Holland and Diamond (2005) while more recent efforts have developed reduced models for pedestal ETG transportHatch et al. (2022) and multiscale quasilinear saturation rules.Staebler et al. (2016) Additionally, a scale-separated model of coupled gyrokinetic equationsHardman et al. (2019) has shown that ion-scale turbulence influences electron-scale dynamics through parallel-to-the-field shearing, which suppresses the ETG growth rate. Hardman, Barnes, and Roach (2020)
The goal of this work is to account for heat losses due to ETG turbulence in global ion-scale simulations and to probe the effects of ETG turbulence on the ITG background. The paper is outlined as follows. Section II describes a theoretical model focused on electron-scale effects in global ion-scale gyrokinetic simulation. Local ITG and ETG simulations are carried out in GENE to test for a valid scale-separated scenario and the results are described in Section III. Section IV then describes the global ITG simulation in GEM which is followed by the inclusion of a kinetic source term that accounts for excess electron-scale thermal losses from ETG turbulence in GEM. Section V concludes with future plans to couple the ETG streamer potential and intermediate-scale zonal flowsColyer et al. (2017); Holland, Howard, and Grierson (2017); Tirkas et al. (2023) found in ETG simulations to the ion scale.
@@ II Subgrid ETG Model
The gyrokinetic framework for modeling microturbulence in tokamak plasmas assumes an expansion in the parameter $\epsilon=\rho/a\ll 1$, where $\rho$ is the species’ gyroradius involved in the generation of instabilities and $a$ is the device minor radius.Frieman and Chen (1982); Lee (1983); Dubin et al. (1983) This in part allows for separating dynamical equations between small-scale fluctuating quantities and background equilibrium quantities. A further subsidiary expansion can be made assuming $\sqrt{m_{e}/m_{i}}\ll 1$ to separate the dynamics of ion-scale (IS) and electron-scale (ES) instabilities, such as ITG and ETG modes.Abel and Cowley (2013); Hardman et al. (2019) Distinct equations can then be used to investigate the effects of coupling between the two scales.
The primary assumption of the subgrid model is to take the electron-scale gyrokinetic equation as stand-alone, i.e. unaffected by ion-scale turbulence. Then electron-scale effects are averaged over intermediate scales in time and space to capture the effects of ETG turbulence on the ITG turbulent background, and effects from ETG flux-tube simulations are added into global ITG simulations. The amplitude of the flux-tube steady-state is varied in accordance with values reported in multiscale simulations involving core ITG and ETG turbulence and the effect of new terms are either discussed further or investigated.
The subgrid model might further be incorporated with the multiscale model of Ref. Hardman et al., 2019 to perform self-consistent coupled simulations of ITG and ETG turbulence. As the theory in Ref. Hardman et al., 2019 makes various assumptions about the electron-scale turbulence, the final equations focus on the effects of ITG turbulence on the electron scale. These assumptions include gyroBohm flux scaling, perpendicular isotropization, and strong scale separation, which are not always seen in local multiscale simulations when ETG turbulence effects are important. The primary goal here is rather to recreate effects of electron-scale turbulence in multiscale simulations by making few assumptions about ES turbulence and adding ES effects directly from flux-tube simulation.
In focusing on the combined effects of ITG and ETG turbulence, the subgrid model retains only electrostatic effects. The governing gyrokinetic Vlasov equation takes the form
@@
where $\langle\dots\rangle_{\alpha}$ represents a gyro-phase angle average. For simplicity, the gyroaverage notation is dropped from now on.
@@
where $\delta f^{\prime}$ retains new components that may be generated in the original IS distribution function due to the inclusion of ES ExB effects. In general this could also include cross-scale energy cascading effects, though this is ignored with the stand-alone ES assumption. An averaging procedure over intermediate mesoscales in time and in space perpendicular to the field, $\tau_{\textrm{m}}$ and $l_{\textrm{m}}$, can then be introduced, where the intermediate values lie between IS and ES scales defined by $a\gg\rho_{i}\gg\rho_{e}$ and $\tau\gg\omega_{*,i}^{-1}\gg\omega_{*,e}^{-1}$. Here, $\tau$ is the transport timescale and $\omega_{*,s}$ are the respective ion and electron diamagnetic frequencies. The choice of intermediate-scale values is discussed further in Section III in reference to the ES flux-tube simulation results.
The mesoscale average of the fluctuations results in retaining only IS fluctuations:
@@
Here, $\langle\dots\rangle_{\textrm{m}}$ represents a mesoscale average in perpendicular space and time, and it is assumed the mesoscale average of $\delta f^{\prime}$ recovers the IS distribution function. As the electron scale is stand-alone, the ES dynamics are described by the gyrokinetic equation
@@
$=-\frac{1}{B}\delta\textbf{E}_{\textrm{ES}}\times\textbf{b}\cdot\nabla f_{0}+q(v_{\parallel}\textbf{b}+\textbf{v}_{D})\cdot\delta\textbf{E}_{\textrm{ES}}\frac{f_{0}}{T}+S_{\textrm{ES}},$
where $S_{\textrm{ES}}$ represents a source term that has been added to ensure a steady-state consistent with the flux-tube approximation. To obtain an equation for the remaining scales, the total fluctuating quantities are substituted into (1) and then the ES equation subtracted, resulting in
@@
Equation (5) is then averaged over the intermediate mesoscale to find the new IS gyrokinetic equation
@@
The three new terms are grouped together on the middle line of (6) for clarity. These new terms represent the averaged effects of the electron-scale turbulence in ion-scale simulations. The first term represents the additional guiding-center motion due to the ETG field. The second term is due to effects of the ITG field on the ETG distribution function, which averages to zero according to (3). The final term, $S_{\textrm{ES}}$, is used to account for electron thermal transport by ES turbulence in IS simulations. Any average effect of collisions at electron-scales is ignored assuming low collision frequency in the core.
Figure 1: (a) Denisty and temperature profiles as a function of $r/a$ and (b) normalized gradient profiles as defined by (9) (10). (c) ITG and ETG growth rates from linear electrostatic GENE simulations at $r/a=0.50$.
A form for the source term can be found by considering flux-surface-averaging (4). Due to the periodic boundary conditions in the flux-tube approximation one finds the source term must come from the ES ExB nonlinearity,
@@
where $\langle\ldots\rangle_{\psi}$ represents a flux-surface average. As this source is responsible for maintaining the steady-state, it can account for excess electron thermal transport from ETG modes. The second non-zero term is recast as a diffusion operator in real space,
@@
As the ion response to ETG turbulence is adiabatic, only the electron distribution function is modified in GEM. Flux-tube ETG simulations are carried out using GENE at different radial locations and a kinetic form of $S_{\textrm{ES}}$ is included in global IS GEM simulations to capture fine-scale electron heat transport. The source added to GEM is varied within reasonable expectations as reported in multiscale simulation.Holland, Howard, and Grierson (2017) Radial theories for the global ES turbulence profile are then compared that allow for using linear simulation results, requiring only one nonlinear ES simulation. Future plans to include ES ExB effects are discussed along with effects from intermediate-scale zonal flows, which would require breaking scale-separation assumptions.
@@ III Local Simulations
with $\hat{s}=\frac{r}{q}\frac{dq}{dr}$ the magnetic shear profile. To retain only electrostatic instabilities of interest, the plasma beta factor, $\beta=8\pi n_{0e}T_{0e}/B_{0}^{2}$, with $B_{0}$ the on-axis magnetic field, is set to $\beta=1e^{-4}$.
Table 1: Time-averaged heat fluxes at $r/a=0.5$ for all simulations, in gyroBohm units.
@@
Figure 2: (a) Electron heat flux $k_{y}$ spectra and electrostatic potential spectra for (b) $k_{y}$ and (c) $k_{x}$ for GENE IS and ES flux-tube scenarios. Labels ‘es’ and ‘em+$\gamma_{E^{\prime}}$’ correspond to electrostatic runs and electromagnetic runs with shear flow. Averaging brackets correspond to spatial averaging over variables which aren’t retained on the $x$-axis from $k_{x}$, $k_{y}$, and $z$.
@@
The normalized mode frequencies, $\omega/(c_{s}/R)$, and growth rates, $\gamma/(c_{s}/R)$, for the instabilities are calculated using GENE linear electrostatic simulations. Here $c_{s}=\sqrt{T_{e}/m_{D}}$ is the deuteron sound speed. The linear growth rates at $r/a=0.5$ are shown in Fig. 1 as a function of $k_{y}\rho_{i}$ which spans from ITG to ETG scales. Grid convergence values in $z\times v_{\parallel}\times\mu$ are found by increasing grid resolution until growth rates are constant to three decimal places. Here, $z$ is the field-line-following coordinate, $v_{\parallel}$ is the particle velocity along a field line, and $\mu=mv_{\perp}^{2}/qB$ is the magnetic moment which represents a particle’s velocity perpendicular to a field line. Convergence was checked for the most unstable mode as well as a longer wavelength mode near the peak of the nonlinear heat flux spectrum.
@@
The resulting grid resolutions, $z\times v_{\parallel}\times\mu$, used for linear simulations were $32\times 48\times 16$ at the ion scale and $48\times 48\times 48$ at the electron scale. For both scales $32$ gridpoints are taken in the local radial coordinate $x$. The IS modes start at $k_{y}\rho_{i}=0.05$ with $40$ modes up to $k_{y}\rho_{i}=2.0$ and the ES modes start at $k_{y}\rho_{i}=2.0$ with $32$ modes up to $k_{y}\rho_{i}=64.0$, where $k_{y}$ is the wavenumber in the binormal, $y$, direction. A clear scale separation in both time and space - $\gamma$ and $k_{y}$ can be seen in Fig. 1 which as expected is on the order of $\sqrt{m_{D}/m_{e}}\sim 60$. Growth rates for both scales were appreciable out to $r/a=\{0.35,0.65\}$ and fell to zero by $r/a=\{0.2,0.8\}$. While the linear simulation results may include trapped-electron modes (TEMs), primarily in the range $k_{y}\rho_{i}\sim 0.5-2$, the nonlinear GENE results presented next show little radial heat flux or electrostatic potential in $k_{y}$ at this intermediate range.
@@
The nonlinear simulations are discussed here in detail for both scales at $r/a=0.5$. The nonlinear $z\times v_{\parallel}\times\mu$ grid resolution was reduced to lower values for which growth rates were still converged to within ±$\pm 1\%$. The new grid resolutions are $16\times 32\times 8$ and $32\times 32\times 16$ for the IS and ES scales. The ITG runs include $32$ modes ranging from $k_{y}\rho_{i}=0.0625$ to $2.0$, and ETG runs include $32$ modes ranging from $k_{y}\rho_{i}=2.0$ to $64.0$. The perpendicular domain sizes, $L_{x}\times L_{y}$, are $96\rho_{i}\times 100\rho_{i}$ and $360\rho_{e}\times 190\rho_{e}$, with the $L_{x}:L_{y}$ ratio increased for ETG runs to allow for possible generation of intermediate-scale zonal flows. The radial grid resolutions are $\Delta x=0.50\rho_{i}$ and $\Delta x=0.62\rho_{e}$ respectively. Ion and electron heat fluxes, $Q_{i}$ and $Q_{e}$, are listed in Table 1 for both the ES (#$\#1$) and IS (#$\#2$) nonlinear runs and compared to the results of the global GEM simulation (#$\#5$) at $r/a=0.50$ which is later described in Section IV. Heat fluxes are normalized using the gyroBohm value $Q_{gB}=n_{e}c_{s}T_{e}(\rho_{D}/R)^{2}$, with $\rho_{D}$ the deuteron gyroradius.
@@
The inclusion of background ExB shear and electromagnetic effects are also considered at both scales to illustrate a more realistic experimental scenario in which ITG modes are more strongly regulated.Kim, Horton, and Dong (1993); Kinsey, Waltz, and Candy (2005) The choice of shearing rate, $\gamma_{E^{\prime}}\approx\gamma_{\textrm{ITG}}^{\max}/2=0.2$, comes from comparing the most unstable ITG growth rate in the electromagnetic case to the maximum shearing rate within experimental uncertainty for the DIII-D IBS values from Ref. Holland, Howard, and Grierson, 2017. The heat fluxes for electromagnetic runs with shear flow are labeled #$\#3$ and #$\#4$ for the ES and IS simulations. These scenarios then provide reasonable and strong levels of ETG turbulence to use in comparison to multiscale simulations results. Specifically, using runs #$\#1$ and #$\#5$ together, the ratio of ES to total electron heat transport falls in the approximate range of $1/2-2/3$ seen in multiscale simulations with appreciable ES effects.Howard et al. (2016); Holland, Howard, and Grierson (2017)
@@
For Runs #$\#1-4$, the spatial averages of the heat flux spectra are shown as a function of $k_{y}$ in Fig. 2. The spatial averages of the electrostatic potential spectra are further shown as functions of $k_{y}$ and $k_{x}$ in Figs. 2 and 2. Electrostatic simulation results are labeled ‘es’, while the electromagnetic results with shear flow are labeled ‘em+$\gamma_{E^{\prime}}$’. Differences in ITG turbulence levels between the ‘es’ and ‘em’ cases is clear. There is also a strong spatial separation of scales in $k_{y}$ for both the heat flux and potential spectra. However, the $k_{x}$ potential spectrum is continuous, due to the generation of intermediate-scale zonal flows at the electron scale. Such zonal flows have been reported in multiscale simulations with weakHolland, Howard, and Grierson (2017) and subcriticalHoward et al. (2016) ITG turbulence. Capturing these effects requires breaking scale-separation assumptions and is discussed further in Section V.

### v3

# A Subgrid Model for Electron-Scale Turbulent Transport in Global Ion-Scale Gyrokinetic Simulations of Tokamak Plasmas

@@ Abstract
A subgrid electron-temperature-gradient (ETG) model is demonstrated here which averages local electron-scale turbulence over intermediate scales in space and time to include in global ion-temperature-gradient (ITG) simulations. This approach results in ion-scale equations which incorporate the electron heat transport from ETG turbulence and effects of electron-scale turbulence on the ion scale. Flux-tube ETG Cyclone Base Case simulations are performed at different radial locations and a kinetic form of the flux is added to global ion-scale simulations as a source term. Analytic radial profiles of ETG heat flux are constructed and compared to flux-tube results at multiple radial locations. Different ratios of ITG to ETG heat flux levels are considered and the results of capturing ETG heat transport in global ion-scale simulations are discussed. Potential coupling of the ETG streamer potential and intermediate-scale zonal flows to the ion scale is further addressed.
@@ I Introduction
The role of ETG turbulence has been studied in various tokamak scenariosGurchenko and Gusakov (2010); Chapman et al. (2015); Ren et al. (2017); Ryter et al. (2019); Kiefer et al. (2021) and is particularly important in cases of marginal ion-scale turbulence.Ren et al. (2024) Specifically for ITER, alpha-particle and electron cyclotron heating effects are expected to drive meaningful ETG turbulence levels, leading to important multiscale dynamics.Holland, Howard, and Grierson (2017); Howard et al. (2017) To better understand interactions between the disparate scales, local multiscale simulations of core ITG and ETG turbulence have garnered much interest.Jenko (2004); Maeyama et al. (2015); Howard et al. (2016); Holland, Howard, and Grierson (2017); Howard et al. (2017); Bonanomi et al. (2018); Howard et al. (2021) In general, turbulent spectra are distinctly scale-separated, and cross-scale interactions lead to changes in steady-state transport levels which can better predict experimental losses for both species. A recent overview of multiscale simulation results can be found in Ref. Maeyama et al., 2024.
As multiscale simulations require resolving electron gyroradius scales, the sizes of simulation domains become limited. Consequently, reduced modeling of electron-scale turbulence is particularly valuable for whole-device modeling efforts in future burning plasma experiments. Previous work has considered the importance of cross-scale interactions,Itoh et al. (2002); Jenko (2004); Holland and Diamond (2005) while more recent efforts have developed reduced models for pedestal ETG transport Hatch et al. (2022) and multiscale quasilinear saturation rules.Staebler et al. (2016) Additionally, a scale-separated model of coupled gyrokinetic equationsHardman et al. (2019) has shown that ion-scale turbulence influences electron-scale dynamics through parallel-to-the-field shearing which suppresses the ETG growth rate.Hardman, Barnes, and Roach (2020)
The goal of this work is to account for heat losses due to ETG turbulence in global ion-scale simulations and to probe the effects of ETG turbulence on the ITG background. The paper is outlined as follows. Section II describes a theoretical model which accounts for electron-scale effects in global ion-scale gyrokinetic simulation. Local ITG and ETG simulations are carried out in GENE to test for a valid scale-separated scenario and the results are described in section III. Section IV then describes the global ITG simulation in GEM. This is followed by the results of including a kinetic source term that accounts for excess electron-scale thermal losses due to ETG turbulence in GEM. Section V concludes with future plans to couple the ETG streamer potential and intermediate-scale zonal flowsColyer et al. (2017); Holland, Howard, and Grierson (2017); Tirkas et al. (2023) found in ETG simulations to the ion scale.
@@ II Subgrid ETG Model
The gyrokinetic framework for modeling microturbulence in tokamak plasmas assumes an expansion in the parameter $\epsilon=\rho/a\ll 1$, where $\rho$ is the species’ gyroradius involved in the generation of instabilities and $a$ is the device minor radius.Frieman and Chen (1982); Lee (1983); Dubin et al. (1983) This in part allows for separating dynamical equations between small-scale fluctuating quantities and background equilibrium quantities. A further subsidiary expansion can be made using the small electron-to-ion mass ratio, $\sqrt{m_{e}/m_{i}}\ll 1$, to separate the dynamics of ion-scale (IS) and electron-scale (ES) instabilities, such as ITG and ETG modes. Abel and Cowley (2013); Hardman et al. (2019) Distinct equations can then be used to investigate the effects of coupling between the two scales. Note that a lowercase ‘es’ is instead used to refer to electrostatic simulations further on.
The primary assumption of the subgrid model is to take the electron-scale gyrokinetic equation as stand-alone, i.e. unaffected by ion-scale turbulence. Then local ES turbulence is averaged over intermediate scales in time and space perpendicular to the field to include effects of ETG turbulence in global ITG simulations. The amplitude of the ES flux-tube steady-state is varied in accordance with values reported in multiscale simulations involving core ITG and ETG turbulence and the effect of new terms are either discussed further or investigated.
The subgrid model might further be incorporated with the multiscale model of Ref. Hardman et al., 2019 to perform self-consistent coupled simulations of ITG and ETG turbulence. As the theory in Ref. Hardman et al., 2019 makes various assumptions about the electron-scale turbulence, the final equations focus on the effects of ITG turbulence on the electron scale. The goal here is to recreate effects of electron-scale turbulence in multiscale simulations by using minimal assumptions and incorporating ES effects directly from flux-tube simulation. The intermediate-scale averaging procedure offers a reduced approach for including ES turbulent effects, making it practical for use in global ion-scale simulations. The accuracy of the subgrid model in capturing cross-scale interactions can be validated by comparing to local IS simulations that directly include ES turbulence.
In focusing on the combined effects of ITG and ETG turbulence, the subgrid model includes only electrostatic effects. The governing gyrokinetic Vlasov equation takes the form
@@
where $\langle\dots\rangle_{\alpha}$ represents a gyro-phase angle average. For simplicity, the gyroaverage notation is dropped from now on. Collisions are ignored assuming low collision frequency in the core.
@@
where $\delta f^{\prime}$ retains new components that may be generated in the original IS distribution function due to the inclusion of ES ExB effects. In general this could also include cross-scale energy cascading effects, though this is ignored with the stand-alone ES assumption. An averaging procedure over intermediate mesoscales in time and in space perpendicular to the field, $\tau_{\textrm{m}}$ and $l_{\textrm{m}}$, can then be introduced, where the intermediate values lie between IS and ES scales defined by $a\gg\rho_{i}\gg\rho_{e}$ and $\tau\gg a/v_{th,i}\gg a/v_{th,e}$. Here, $\tau$ is the transport timescale and $v_{th}$ is the particle thermal velocity. The specific choice of intermediate-scale values is described further in section III in reference to the ES flux-tube simulation results.
Mesoscale averaging the components of (2) results in retaining only IS fluctuations:
@@
Here, $\langle\dots\rangle_{\textrm{m}}$ represents a mesoscale average in space perpendicular to the field and time. It is assumed the mesoscale average of $\delta f^{\prime}$ recovers the IS distribution function. As the electron scale is stand-alone, the ES dynamics are described by the gyrokinetic equation
@@
$=-\frac{1}{B}\delta\textbf{E}_{\textrm{ES}}\times\textbf{b}\cdot\nabla f_{0}+q(v_{\parallel}\textbf{b}+\textbf{v}_{D})\cdot\delta\textbf{E}_{\textrm{ES}}\frac{f_{0}}{T}+S_{\textrm{ES}}.$
A source term, $S_{\textrm{ES}}$, has been explicitly included to ensure a steady-state consistent with the flux-tube approximation. To obtain an equation for the remaining scales, the total fluctuating quantities are substituted into (1) and then the ES equation subtracted, resulting in
@@
Equation (5) is then averaged over the intermediate mesoscales to find the new IS gyrokinetic equation
@@
The three new terms are grouped together on the middle line of (6) for clarity. These new terms represent the averaged effects of electron-scale turbulence in ion-scale simulations. The first term represents the additional guiding-center motion due to the ETG field. The second term is due to effects of the ITG field on the ETG distribution function, which averages to zero according to (3). The final term, $S_{\textrm{ES}}$, is used to account for electron thermal transport by ES turbulence in IS simulations.
Figure 1: (a) Density and temperature profiles and (b) the respective normalized gradient profiles as functions of $r/a$ defined by (9) and (10). (c) ITG and ETG growth rates from linear electrostatic GENE simulations at $r/a=0.50$.
Given local electron-scale simulations, a form for the source term can be found by taking the flux-surface-average of (4). Due to the periodic boundary conditions in the flux-tube approximation one finds that the source term must come from the ES ExB nonlinearity,
@@
where $\langle\ldots\rangle_{\psi}$ represents a flux-surface average. Note that the mesoscale average can be used if the turbulence is strongly scale separated, but in principle the linear terms may not average to zero for intermediate-scale modes. As this source is responsible for maintaining the steady-state transport, it can account for excess electron thermal transport from ETG modes.
The second non-zero term is recast as a diffusion operator in real space,
@@
Sections III and IV address including the ES source term in global ion-scale simulations to capture electron-scale heat transport. Flux-tube ETG simulations are carried out using GENE at different radial locations. A kinetic source term is constructed using local results at the peak temperature gradient location and the source term is added to global IS GEM simulations. The ES source term is varied in accordance with various multiscale simulation results, and the result of adding ES transport to global IS simulation is discussed. Different analytic radial profiles of the ETG heat flux are then compared to results from local ES simulations at multiple radial locations.
@@ III Local Simulations
The density and temperature profiles and their gradients are shown in Fig. 1 and 1. The safety factor profile is given by
@@
with $\hat{s}=\frac{r}{q}\frac{dq}{dr}$ the magnetic shear profile. To retain only the electrostatic instabilities of interest, the plasma beta factor, $\beta=8\pi n_{0e}T_{0e}/B_{0}^{2}$, with $B_{0}$ the on-axis magnetic field, is taken to be $\beta=1e^{-4}$. Multiscale CBC simulations have previously reported important cross-scale interactions,Maeyama et al. (2015) and so the simplified circular CBC parameters were chosen to facilitate initial theoretical investigation. While these parameters are idealized, ETG turbulence profiles can in principle be captured for any experimental scenario by using results from multiple flux-tube simulations at different radial locations.
Table 1: Time-averaged heat fluxes for all simulations, in gyroBohm units.
@@
Figure 2: Flux spectra for GENE IS and ES flux-tube scenarios. (a) Electron heat flux and (b) electrostatic potential spectra as functions of $k_{y}$, averaged over $k_{x}$ and $z$. (c) Electrostatic potential spectra as a function of $k_{x}$, averaged over $k_{y}$ and $z$. Labels ‘es’ and ‘em+$\gamma_{E^{\prime}}$’ correspond to electrostatic runs and electromagnetic runs with shear flow respectively.
@@
The normalized mode frequencies, $\omega/(c_{s}/R)$, and growth rates, $\gamma/(c_{s}/R)$, for the instabilities are calculated using GENE linear electrostatic simulations. Here $c_{s}=\sqrt{T_{e}/m_{D}}$ is the deuteron sound speed. The linear growth rates at $r/a=0.50$ are shown in Fig. 1 spanning IS and ES scales in $k_{y}\rho_{i}$, where $k_{y}$ is the wavenumber in the binormal direction. Grid convergence values in $z\times v_{\parallel}\times\mu$ were found by increasing grid resolution until growth rates were constant to three decimal places. Here, $z$ is the field-line-following coordinate, $v_{\parallel}$ is the particle velocity along a field line, and $\mu=mv_{\perp}^{2}/2B$ is the magnetic moment which represents particle velocity perpendicular to a field line. Convergence was checked for the most unstable mode as well as a longer wavelength mode closest to the peak of the nonlinear flux spectra.
@@
The resulting grid resolutions, $z\times v_{\parallel}\times\mu$, used for linear simulations were $32\times 48\times 16$ at the ion scale and $48\times 48\times 48$ at the electron scale. For both scales $32$ gridpoints are taken in the local radial coordinate $x$. The IS modes start at $k_{y}\rho_{i}=0.05$ with $40$ modes up to $k_{y}\rho_{i}=2.0$ and the ES modes start at $k_{y}\rho_{i}=2.0$ with $32$ modes up to $k_{y}\rho_{i}=64.0$. A clear scale separation in both frequency and wavenumber can be seen in Fig. 1. The scale separation is of order $\gamma_{ES}/\gamma_{IS}\sim k_{y,ES}/k_{y,IS}\sim\sqrt{m_{D}/m_{e}}\sim 60$, which is expected given the respective IS and ES orderings of $\rho_{i}$ and $\rho_{e}$ in space perpendicular to the field and $a/v_{th,i}$ and $a/v_{th,e}$ in time. With the equilibrium gradients dropping off outward from the center, growth rates for both scales remained appreciable out to $r/a=0.35$ and $0.65$ while falling to zero near $r/a=0.20$ and $0.80$. While the linear simulation results may include trapped-electron modes (TEMs), primarily in the range $k_{y}\rho_{i}\sim 0.50-2.0$, the nonlinear GENE results presented next show little radial heat flux or electrostatic potential in $k_{y}$ at this intermediate range.
@@
The nonlinear simulations are discussed here in detail for both scales at $r/a=0.50$. The nonlinear $z\times v_{\parallel}\times\mu$ grid resolution was reduced to lower values for which growth rates were still converged to within ±$\pm 1\%$. The new grid resolutions are $16\times 32\times 8$ and $32\times 32\times 16$ for the IS and ES scales. Nonlinear IS simulations were also tested with an increased resolution of $32\times 48\times 16$, while for ES simulations the resolution in each of the three dimensions was increased by a factor of $2$ respectively to test for convergence. The ITG runs include $32$ modes ranging from $k_{y}\rho_{i}=0.0625$ to $2.0$, and ETG runs include $32$ modes ranging from $k_{y}\rho_{i}=2.0$ to $64.0$. The perpendicular domain sizes, $L_{x}\times L_{y}$, are $96\rho_{i}\times 100\rho_{i}$ and $360\rho_{e}\times 190\rho_{e}$, with the $L_{x}:L_{y}$ ratio increased for ETG runs to allow for possible generation of intermediate-scale zonal flows. The radial grid resolutions are $\Delta x=0.50\rho_{i}$ and $\Delta x=0.62\rho_{e}$ respectively. The perpendicular domain for ETG runs was tested by separately increasing $L_{x}$ and $L_{y}$ by a factor of $2$ each.
@@
Ion and electron heat fluxes, $Q_{i}$ and $Q_{e}$, are listed in Table 1 for all ES and IS nonlinear runs. Heat fluxes are normalized using the gyroBohm value $Q_{gB}=n_{e}c_{s}T_{e}(\rho_{D}/R)^{2}$, with $\rho_{D}$ the deuteron gyroradius. Runs #$\#1$ and #$\#2$ correspond to ES and IS simulations without background shear flow or electromagnetic effects. Runs #$\#3$ (ES) and #$\#4$ (IS) include these effects to regulate ITG turbulenceKim, Horton, and Dong (1993); Kinsey, Waltz, and Candy (2005) in the unstable CBC scenario, providing a comparison to multiscale scenarios in which ITG turbulence is marginal. This results in a $95\%$ reduction in $Q_{i}$. By comparison, simulations with only shear flow or finite-beta effects show reductions of approximately $37.5\%$ and $85\%$ respectively. The value of $\beta$ is derived directly from the equilibrium profiles. The choice of shearing rate, $\gamma_{E^{\prime}}\approx\gamma_{\textrm{ITG}}^{\max}/2=0.2$, comes from comparing the most unstable ITG growth rate in the electromagnetic case to the maximum shearing rate within experimental uncertainty for the DIII-D IBS values from Ref. Holland, Howard, and Grierson, 2017. The heat fluxes are further compared to results from the global ion-scale GEM simulation (run #$\#5$) which is discussed further in section IV. The GEM simulation ranges from $r/a=0.20$ to $0.80$, and the heat flux is calculated by averaging over a centered domain spanning from $r/a=0.35$ to $0.65$.
These scenarios provide variable ratios of IS and ES turbulence for comparison with multiscale simulations results. In runs #$\#1$ and #$\#5$, the ratio of ES to total electron heat transport falls within the approximate range of $1/2-2/3$ seen in multiscale scenarios with appreciable ES effects.Howard et al. (2016); Holland, Howard, and Grierson (2017) Finite-beta and shear flow effects are not included directly in GEM as the CBC scenario with electromagnetic effects is often numerically challenging.Waltz (2010) Instead, ES effects are magnified by a factor of $3$ when included in GEM in section IV. This increased factor is used to match the ratio of electron heat flux in runs #$\#3$ and #$\#4$ when ITG is regulated. The magnified ES scenario represents cases of marginal ITG turbulence in multiscale scenarios, where the ratio of total $Q_{e}:Q_{i}$ can range from $1$-$3$.Howard et al. (2016); Holland, Howard, and Grierson (2017); Bonanomi et al. (2018)
@@
For Runs #$\#1-4$, the spatial averages of the heat flux spectra are shown as functions of $k_{y}$ in Fig. 2. The spatial averages of the electrostatic potential spectra are further shown as functions of $k_{y}$ in Fig. 2 and $k_{x}$ in Fig. 2. Electrostatic simulation results are labeled ‘es’, while the electromagnetic results with shear flow are labeled ‘em+$\gamma_{E^{\prime}}$’. The difference in ITG turbulence levels between both cases is clear. A strong spatial separation of scales exists in $k_{y}$ for both the heat flux and potential spectra. However, the $k_{x}$ potential spectrum is continuous, due to the generation of intermediate-scale zonal flows at the electron scale. Such zonal flows have been reported in multiscale simulations with marginal levels of ITG turbulence, where they may further suppress ion-scale fluctuations.Holland, Howard, and Grierson (2017) Capturing these effects would require breaking scale-separation assumptions and this is briefly mentioned further in section V.